13. Use the appropriate normal distribution to approximate the resulting binomial distribution. Bearings is the principal supplier of ball bearings for the Sperry Gyroscope Company. It has been determined that 6% of the ball bearings shipped are rejected because they fail to meet tolerance requirements. What is the probability that a shipment of 200 ball bearings contains more than 14 rejects?
Answer #1

Answer

sample size is large enough because we have 200 ball bearings.

probability p = 6/100 = 0.06

So, probability is niether near to 0 nor near to 1

so, we can use the normal approximation.

where mean = np and σ=Vn * p * (1-p)

we have n = 200 and p = 0.06

setting the values, we get

mean = 200*0.06 = 12

and standard deviation = σ = V/200 * 0.06 * (1-0.06) = V 1 1.28 3.36

Now, we have to find the probability that more than 14 ball bearings rejected

So, P(x>14) = P(z>(ar{x}-mu)/sigma)

where ar{x} = 14, n=200, mu=12, sigma=3.36

setting the given values, we get

P14) Pz> (14 - 12)/3.36)- P(z >(2/3.36))-P(z >0.5952)

using identity P(z>a) = 1-P(z<a)

wecan write it as

P(z > 0.5952) = 1-P(z < 0.5952) 1-0.7257 = 0.2743 (using z distribution table)

So, the required probability is 0.2743